MAGNETIC ROTATION STUDY OF THE A^{3}Π_{1u}–X^{1}Σ_{g}^{+} SYSTEM OF ^{79}Br_{2}

\underline{C.D. BOONE}, A. CHANDA^{a}, F.W. DALBY, and I. OZIER, Department of Physics and Astronomy, University of British Columbia, 6224 Agricultural Road, Vancouver, BC, Canada, V6T 1Z1.

The magnetic rotation spectrum of ^{79}Br_{2} has been measured from 14,988 to 15,488 cm^{−1} using a tunable, cw dye laser (effective linewidth ∼ 2 MHz) with DCM dye. The study was carried out using a cell filled to ∼6.5 Torr with 99.4% isotopically pure ^{79}Br_{2}. Over 5000 transitions (with Doppler width of about 450 MHz) were measured, the majority of which (> 4500) belong to the A^{3}Π_{1u}–X^{1}Σ_{g}^{+} system of the molecule. The range of upper-state vibrational levels observed (v′ = 13 to 38) extended very close to the dissociation limit of the A state. With more standard techniques such as absorption or fluorescence, transitions in the A–X system are typically difficult to detect because of much stronger transitions in the same region from the B^{3}Π_{0^{+}u}–X^{1}Σ_{g}^{+} system; the opposite is true for magnetic rotation, due to magnetic activity of the A^{3}Π_{1u} electronic state. Magnetic rotation signals normally get weaker with increasing J, since the effective g-factor, g_{J}, decreases as J increases. As a result, magnetic rotation spectra usually give strong signals only for low-J transitions. For the A–X system of Br_{2}, however, an interference effect between spin-uncoupling and magnetic mixing to other electronic states—primarily the A^{′3}Π_{2u} state—led to the measurement of rotational transitions right up to the breaking-off point (e.g. J = 82 for v′ = 25). These high-J transitions were comparable in strength with those at low-J. The interference effect manifests itself as the addition to g_{J} of a term that increases with J. Structure observed in low-J transitions has been attributed to hyperfine effects in the A state. Transitions to quasibound levels above the dissociation limit of the A state appear to suffer from strong, systematic perturbations, likely due to interaction with a dissociative state. Perhaps the most interesting feature of the spectrum is a series of extra lines tentatively identified as belonging to the A^{′3}Π_{2u}–X^{1}Σ_{g}^{+} system, the result of intensity-borrowing from near-resonant perturbations with the A state v′ = 27 rotational levels 2 ≤ J ≤ 31.

a Present address: Unisearch Associates Inc., 96 Bradwick Drive, Concord, ON, Canada, L4K 1K8.